OpenStudy (anonymous):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bars sell for $5 per pound and the milk chocolate fudge bars sell for $4 per pound. You purchased a combination of these bars for your family and friends and your bill came to $150. This scenario can be modeled by the equation 5d + 4m = 150. Which of the following is not a possible combination of dark chocolate fudge bars and milk chocolate fudge bars? 6 dark bars and 30 milk bars 10 dark bars and 25 milk bars 12 dark bars and 23 milk bars 14 dark bars and 20 milk bars

OpenStudy (campbell_st):

the easy way to solve this is by substitution each combination 6 dark 30 milk 5 x 6 + 30 x 4 = 150 I'll let you do the rest

OpenStudy (ash2326):

We have \[5d+4m=150\] d=no. of dark chocolates m= no. of milk chocolate Now we have different options, substitute each and check which one satisfies I'll show you one, the rest you try I'll solve for 2nd option=10 dark and 25 milk so d=10 and m=25 \[5 \times 10+4 \times 25=150\] \[50+100=150\] This satisfies, so this one of the correct options, you check the rest @AngelsUndoing12